A small 2.00 g plastic ball is suspended by a 20.0 cm long string in a uniform electric field. If the ball is in equilibrium when the string makes a 15.0 degree angle with the vertical, what is the net charge on the ball?
1
Expert's answer
2020-11-06T10:18:20-0500

Let’s draw a free-body diagram:



Here, "W=mg" is the force of gravity (or weight of the plastic ball), "F_e=qE" is the electric force that acts on the plastic ball, "T" is the force of tension in the string, "T_x","T_y" is the projections of the force of tension on axis "x" and "y", respectively, "\theta" is the angle of inclination of the string to the vertical, "q" is the charge of the plastic ball and "E" is the intensity of the electric field.

Let’s write the conditions of the equilibrium for the plastic ball:


"\sum F_x=0, \sum F_y=0."


Let’s consider the forces that act on the plastic ball in the horizontal "x"- and vertical "y"-direction:


"\sum F_x=F_e-Tsin\theta=qE-Tsin\theta=0, (1)""\sum F_y=Tcos\theta-mg=0. (2)"

We can express the force "T" from the second equation:


"T=\dfrac{mg}{cos\theta}."

Then, we can substitute it into the first equation and get:


"qE-mgtan\theta=0,""qE=mgtan\theta."

From the last equation we can find the magnitude of the net charge on the ball:


"q=\dfrac{mgtan\theta}{E}."

Unfortunatelly, we don't know the intensity of the electric field. Let's suppose, for example, that "E=1000\ \dfrac{N}{C}." Then, we can calculate the magnitude of the net charge on the ball:


"q=\dfrac{0.002\ kg\cdot 9.8\ \dfrac{m}{s^2}\cdot tan15^{\circ}}{1000\ \dfrac{N}{C}}=5.25\cdot 10^{-6}\ C"

Answer:

"q=5.25\cdot 10^{-6}\ C"